# Cream: Consistency Regularized Self-Rewarding Language Models

### v2

@@ 3.3 Proposed Algorithm
Equipped with the above two-stage optimization and the consistency-regularized self-rewarding, we are ready to present the implementation of CREAM in Algorithm 1. The whole framework of CREAM is also illustrated in Figure 2. The algorithm starts from the SFT training to obtain the first model parameter $\bm{\theta}_{1}$ in Line 2. A similar approach is applied in Yuan et al. (2024) to avoid calculating the $\mathcal{L}_{\text{SFT}}$ in the future optimization steps. Then for each $\mathbf{x}_{j}$ in the unlabeled prompt set $\mathcal{D}_{\text{U}}$, $N$ response candidates $\{\mathbf{y}_{i}\}_{i=1}^{N}$ are sampled in Line 5. Then reward scores of these $N$ candidates can be calculated according to Rafailov et al. (2023) by
@@
where we use the initial model parameter $\bm{\theta}_{0}$ as the reference policy $\pi_{\text{ref}}$. Since $\beta\geq 0$ and $\log Z(\mathbf{x}_{j})$ is a constant across different response $\mathbf{y}_{i}$ for the same input prompt $\mathbf{x}_{j}$, we can drop these factors and calculate rewards in Line 6. Specifically, when $t=1$, the rank $K_{ij}$ is calculated based on the reference policy $\bm{\theta}_{0}$ itself. Thus we instead use the likelihood $r_{ij}=\log\pi_{\bm{\theta}_{0}}(\mathbf{y}_{ij}|\mathbf{x}_{j})$ as the reward for this edge case. The rank for these $N$ candidates are therefore obtained in Line 7, where $J_{ij}$ means response $\mathbf{y}_{ij}$ is in the $J_{ij}$-th best in the preference list of $\mathbf{x}_{j}$.
@@
As discussed in Eq. (3.8), a baseline reward model is required to measure the consistency. In the self-rewarding scenario, it is infeasible to add an external reward model as the baseline reward model. Fortunately, we can employ the model before last update $\bm{\theta}_{t-1}$ as the baseline reward model $\bm{\theta}_{t}^{\prime}$ (i.e., last iteration’s model) for evaluating the consistency of the model $\bm{\theta}$, thanks to chances provided by iterative training manner. Such a procedure helps mitigate the training error introduced in $t-1$-th step before obtaining $\bm{\theta}_{t}$. Considering a pair of tied preference pair $\mathbf{y},\mathbf{y}^{\prime}$ both performing well, as demonstrated in Figure 1. $P[\mathbf{y}\succ\mathbf{y}^{\prime}|\mathbf{x},\bm{\theta}_{t}]$ will be oscillating around $0.5$ when $t$ grows due to the random noise. Otherwise $P[\mathbf{y}\succ\mathbf{y}^{\prime}|\mathbf{x},\bm{\theta}_{t}]$ might consistently converge to 00 or $1$. Due to this oscillation, the consistency between $\bm{\theta}_{t-1}$ and $\bm{\theta}_{t}$ on this specific preference pair would be low, and the algorithm will learn less from this noninformative preference pair thus stabilize this oscillation.
@@
Kendall’s Tau coefficient is a widely used coefficient (McLeod, 2005; Abdi, 2007) to measure the consistency of two ranking sequences. Basically, when two sequences perfectly aligns, $\tau_{j}=1$ and when two sequence never aligns, $\tau_{j}=-1$. The following lemma draws a further connection between Kendall’s Tau and the regularization parameter $\lambda$ proposed in Section 3.2.
@@
Given Lemma 3.4, we can recover $\mathcal{C}_{\lambda}\approx 1-\lambda=(1+\tau_{j})/2$ and we use average all $\tau_{j}$ for all $\mathbf{x}_{j}\in\mathcal{D}_{\text{U}}$ in Line 11. Finally, in Line 12, we compose the preference dataset by selecting the best response $\mathbf{y}^{+}_{j}=\mathbf{y}_{i^{+}j}$ and the worst response $\mathbf{y}^{-}_{j}=\mathbf{y}_{i^{-}j}$ which is similar with (Yuan et al., 2024).
$\mathcal{D}_{\text{DPO}}=\{(\mathbf{x}_{j},\mathbf{y}_{i^{+}j},\mathbf{y}_{i^{-}j})|\mathbf{x}_{j}\in\mathcal{D}_{\text{U}},i^{+}=\operatorname*{arg\,min}_{i}J_{ij},i^{-}=\operatorname*{arg\,min}_{i}J_{ij}\}$ (3.11)
@@
$\mathcal{D}_{\text{RDPO}}=\{(\mathbf{x}_{j},\mathbf{y}_{i^{-}j},\mathbf{y}_{i^{+}j})|\mathbf{x}_{j}\in\mathcal{D}_{\text{U}},i^{+}=\operatorname*{arg\,min}_{i}J_{ij},i^{-}=\operatorname*{arg\,min}_{i}J_{ij}\}$ (3.12)
and update $\bm{\theta}_{t+1}$ by minimizing the empirical loss of Eq. (3.7) in Line 14. The detailed proof of theorems and lemmas are provided in the Appendix A.
@@ 4.1 Experimental Setup
Data. In our experiments, we use Open Assistant dataset (Köpf et al., 2024) and only reserve about 3.4K human-annotated examples as the seed SFT data $\mathcal{D}_{\text{S}}$. To construct the unlabeled prompt dataset $\mathcal{D}_{\text{U}}$, we mix prompts of $\mathcal{D}_{\text{S}}$ with the train split of each downstream task (only reserve the prompts) including (1) ARC-Easy/Challenge (Clark et al., 2018), a dataset of genuine grade-school level, multiple-choice science questions-answering dataset; (2) OpenBookQA (Mihaylov et al., 2018), a dataset of elementary-level science questions designed to assess understanding of realistic scenario core facts; (3) SIQA (Sap et al., 2019), a question-answering benchmark for testing models’ abilities to reason about the social implications of everyday events and situations; (4) GSM8K (Cobbe et al., 2021), a primary school level mathematical dataset in which the questions take between 2 and 8 steps to solve. Finally, this process results in a total of 21K prompts in $\mathcal{D}_{\text{U}}$, which we distribute equally across iterative self-rewarding trainings.
@@
Models. Due to limited computational resources, we mainly conduct experiments with two LLMs with about 7B parameters, including Llama-3 (Dubey et al., 2024) and Llama-2 (Touvron et al., 2023), both of which are widely used. Note that the proposed framework is designed for LLMs with any sizes, while validating our findings on other LLMs will be our future work.
@@
Baseline Methods. To validate our findings, we mainly compare our method with SRLM (Yuan et al., 2024) which uses the same LLM to serve as both the policy and reward model to generate preference data for iterative training. Additionally, we introduce a variant of RL with AI feedback (Guo et al., 2024a), referred to as “Oracle”. In this variant, the reward model in SRLM is replaced with an external reward model to demonstrate the upper bound performance of SRLM. Specifically, we use InternLM2 (Cai et al., 2024), a specialized trained reward model, to provide the reward scores for the generated responses. We further enhance Oracle’s rewarding by leveraging the labels from downstream tasks to improve the rewarding accuracy.
@@
Implementation Details. In our experiments, we fine-tune the initial model (M0) on the seed SFT data for $3$ epochs with a learning rate of $61e-$ 6, resulting in model M1. Following SRLM approach, we then iteratively fine-tune the model using DPO in two additional iterations, producing models M2 and M3. In the preference training of each iteration, we set $\beta=0.1$ of DPO loss, and fine-tune the model for $1$ epoch with a learning rate of $61e-$ 6. All training processes use the AdamW optimizer (Loshchilov & Hutter, 2019) with a warmup ratio of $0.1$. For the response sampling stage of all SRLM methods, we use a decoding temperature of $0.8$. For evaluating downstream tasks, we use greedy decoding. All experiments are conducted primarily on NVIDIA A6000 GPUs.
@@ 4.2 Main Results
Table 1: Main results of each method on test sets of downstream tasks. The exact match accuracies are reported. The “” and “” indicate the performance improvement and degradation compared to the method’s last iteration (e.g., M1 $\rightarrow$ M2 and M2 $\rightarrow$ M3), respectively. The best performance among methods using self-rewarding is highlighted in bold.
Model Method Reward Iteration Arc-Easy Arc-Challenge OpenBookQA SIQA GSM8K
Llama-3 Initial - M0 86.29 80.37 86.00 68.58 78.01
SFT - M1 86.78 80.14 86.40 69.50 78.39
Oracle External M2 89.60 82.17 90.00 72.88 80.82
M3 89.31 81.31 90.20 73.75 76.04
SRLM Self M2 87.79 80.38 87.80 70.95 78.01
M3 87.17 81.23 87.30 70.37 77.48
CREAM Self M2 88.89 80.89 88.00 69.79 81.04
M3 89.52 83.36 90.20 72.06 81.73
Llama-2 Initial - M0 61.07 48.98 62.20 50.36 23.65
SFT - M1 60.44 48.46 63.20 50.77 23.88
Oracle External M2 70.20 55.03 75.40 63.66 30.02
M3 71.72 55.80 77.20 62.44 29.57
SRLM Self M2 58.67 46.67 59.80 49.69 25.17
M3 46.55 34.47 49.20 48.06 22.14
CREAM Self M2 58.97 47.53 62.80 50.43 24.41
M3 62.08 48.81 64.60 51.22 25.85
The main results are shown in Table 1. From these results, we observe the following: (1) The Standard SRLM fails to achieve satisfactory performance, particularly with Llama2 which has relatively weaker foundation performance even after SFT fine-tuning (M0 $\rightarrow$ M1), which indicates its limitations for 7B-level LLMs. (2) Compared to SRLM, CREAM achieves a significant improvement across almost all downstream tasks, showing the advantage of introducing the proposed regularization method. (3) SRLM equipped with an external oracle reward model (“Oracle”) can ensure high rewarding accuracy for annotations of self-generated preference data, thereby achieving the best performance overall. This highlights a critical challenge for all methods with unreliable rewarding. Notably, for Llama3, CREAM even outperforms Oracle except on SIQA dataset, showcasing the superior performance of CREAM. This superiority underlines the success of the proposed regularization in mitigating the rewarding bias issue.
Figure 3: Arena of CREAM v.s. SRLM and Oracle with Llama3, judged by GPT-4o.
(4) The consistent performance improvements of CREAM across iterations (M1 $\rightarrow$ M2, M2 $\rightarrow$ M3) validate the effectiveness of our regularization in mitigating the rewarding bias issue. This also demonstrates the benefits of introducing such internal ranking consistency as a regularization signal.
Alignment Arena. To more directly compare the alignment performance of CREAM, we further show the Arena win-rate of our method in Figure 3. We can find that CREAM can beat baseline methods with the same iteration, which confirms the advantage of the proposed regularized self-rewarding. Also, the win and tie rates of CREAM increase with the iteration against stronger model Oracle, indicating the consistent performance improvements across iterations.
@@ 4.3 Analysis
In this section, we provide a detailed analysis to validate and understand the contribution of CREAM in improving the performance via iterative preference fine-tuning manner.
@@ 4.3.1 Analysis of Rewarding
Rewarding Consistency. We first examine the rewarding consistency of different methods using their corresponding models from the last iteration in Table 2. Here, we use the proposed Consistency Rate $\mathcal{C}$, Kendall correlation coefficient $\tau$, Spearman correlation coefficient, and TopOrder metrics to measure the consistency, where the TopOrder metric evaluates whether the final paired preference data remains the same, calculated as follows:
$\text{TopOrder}_{j}=\operatorname{\mathds{1}}\left[\operatorname*{arg\,min}J_{j}=\operatorname*{arg\,min}K_{j}\right]\cdot\operatorname{\mathds{1}}\left[\operatorname*{arg\,max}J_{j}=\operatorname*{arg\,max}K_{j}\right],$
where $J_{j}$ and $K_{j}$ are the rankings of the responses provided by the current model and the last iteration’s model, respectively. This metric assesses whether both the least preferred and most preferred responses are consistently ranked across iterations. The results confirm that SRLMs exhibit a rewarding consistency issue, indicating that the generated preference data may not be reliable for training. In contrast, CREAM can keep the ranking consistency across iterations thanks to the explicit regularization in the training.
Table 2: Ranking consistency of CREAM and SRLM across iterations using Llama3.
@@
Prompt Rewarding v.s. DPO Rewarding. As aforementioned, 7B-level LLMs struggle with generating accurate rewards when using LLM-as-a-Judge prompting (Zheng et al., 2023) due to their limited capacities. We validate this finding in Figure 4 and Figure 5 which compare the ranking accuracy and model performance with prompt rewarding (P-SRLM) against DPO rewarding (SRLM and CREAM) using Llama3, respectively. The ranking accuracy results include the performance on a curated in-domain Preference Data and out-of-domain RewardBench (Lambert et al., 2024) dataset, both of which are formulated as a ranking task to predict the preferred one between two responses. The results suggest that prompt rewarding is not effective for smaller LLMs, as the model performance starts to decrease at the first iteration when trained on the self-rewarded preference data, though the ranking accuracy raises slightly. In contrast, DPO rewarding methods, particularly CREAM, show better performance in both ranking accuracy and five downstream tasks. This is because DPO rewarding is intrinsically aligned with the model’s learning objective.
Ranking Accuracy. The ranking accuracy is crucial for self-rewarding methods, as it directly affects the quality of the preference data for training. Figure 4 directly compares the ranking accuracy of different methods. The Preference Data is curated by collecting self-generated preference pairs, followed by using GPT-4o to annotate the ranking labels. This dataset is used to evaluate the model’s in-domain performance. The RewardBench dataset is used to assess the generalizability of the models beyond the training domain. CREAM consistently achieves higher ranking accuracy than baseline methods, which promises more reliable preference data for training. Even though the overall ranking accuracy ($\leq 70\%$) is not satisfying, the introduced consistency regularization can help mitigate the negative impacts of potentially noisy ranking annotations during preference fine-tuning.

### v3

@@ 3.3 Proposed Algorithm
Equipped with the above two-stage optimization and the consistency-regularized self-rewarding, we are ready to present the implementation of CREAM in Algorithm 1. The whole framework of CREAM is also illustrated in Figure 2. The algorithm starts from the SFT training to obtain the first model parameter $\bm{\theta}_{1}$ in Line 2. A similar approach is applied in Yuan et al. (2024) for avoid calculating the $\mathcal{L}_{\text{SFT}}$ in the future optimization steps. Then for each $\mathbf{x}_{j}$ in the unlabeled prompt set $\mathcal{D}_{\text{U}}$, $N$ response candidates $\{\mathbf{y}_{i}\}_{i=1}^{N}$ are sampled in Line 5. Then reward scores of these $N$ candidates can be calculated according to Rafailov et al. (2023) by
@@
where we use the initial model parameter $\bm{\theta}_{0}$ as the reference policy $\pi_{\text{ref}}$. Since $\beta\geq 0$ and $\log Z(\mathbf{x}_{j})$ is a constant across different response $\mathbf{y}_{i}$ for the same input prompt $\mathbf{x}_{j}$, we can drop these factors and calculate rewards in Line 6. Specially, when $t=1$, the rank $K_{ij}$ is calculated based on the reference policy $\bm{\theta}_{0}$ itself. Thus we instead use the likelihood $r_{ij}=\log\pi_{\bm{\theta}_{0}}(\mathbf{y}_{ij}|\mathbf{x}_{j})$ as the reward for this edge case. The rank for these $N$ candidates are therefore obtained in Line 7, where $J_{ij}$ means response $\mathbf{y}_{ij}$ is in the $J_{ij}$-th best in the preference list of $\mathbf{x}_{j}$.
@@
As discussed in Eq. (3.8), a baseline model is required to measure the consistency. In the self-rewarding scenario, it is infeasible to add an external reward model as the baseline model. Fortunately, we can employ the model before last update $\bm{\theta}_{t-1}$ as the baseline model $\bm{\theta}_{t}^{\prime}$ (i.e., last iteration’s model) for evaluating the consistency of the model $\bm{\theta}$, thanks to chances provided by iterative training manner. Such a procedure helps mitigate the training error introduced in $t-1$-th step before obtaining $\bm{\theta}_{t}$. Considering a pair of tied preference pair $\mathbf{y},\mathbf{y}^{\prime}$ both performing well, as demonstrated in Figure 1. $P[\mathbf{y}\succ\mathbf{y}^{\prime}|\mathbf{x},\bm{\theta}_{t}]$ will be oscillating around $0.5$ when $t$ grows due to the random noise. Otherwise $P[\mathbf{y}\succ\mathbf{y}^{\prime}|\mathbf{x},\bm{\theta}_{t}]$ might consistently converge to 00 or $1$. Due to this oscillation, the consistency between $\bm{\theta}_{t-1}$ and $\bm{\theta}_{t}$ on this specific preference pair would be low, and the algorithm will learn less from this noninformative preference pair thus stabilize this oscillation.
@@
Kendall’s Tau coefficient is a widely used coefficient (McLeod, 2005; Abdi, 2007) to measure the consistency of two ranking sequences. Basically, when two sequences perfectly aligns, $\tau_{j}=1$ and when two sequence never aligns, $\tau_{j}=-1$. The following lemma draws the further connection between the Kendall’s Tau and the regularization parameter $\lambda$ proposed in Section 3.2.
@@
Given Lemma 3.4, we can recover $\mathcal{C}_{\lambda}\approx 1-\lambda=(1+\tau_{j})/2$ and we average all $\tau_{j}$ for all $\mathbf{x}_{j}\in\mathcal{D}_{\text{U}}$ in Line 11. Finally, in Line 12, we compose the preference dataset by selecting the best response $\mathbf{y}^{+}_{j}=\mathbf{y}_{i^{+}j}$ and the worst response $\mathbf{y}^{-}_{j}=\mathbf{y}_{i^{-}j}$ which is similar with (Yuan et al., 2024).
$\mathcal{D}_{\text{DPO}}=\{(\mathbf{x}_{j},\mathbf{y}_{i^{+}j},\mathbf{y}_{i^{-}j})|\mathbf{x}_{j}\in\mathcal{D}_{\text{U}},i^{+}=\operatorname*{arg\,min}_{i}J_{ij},i^{-}=\operatorname*{arg\,max}_{i}J_{ij}\}$ (3.11)
@@
$\mathcal{D}_{\text{RDPO}}=\{(\mathbf{x}_{j},\mathbf{y}_{i^{-}j},\mathbf{y}_{i^{+}j})|\mathbf{x}_{j}\in\mathcal{D}_{\text{U}},i^{+}=\operatorname*{arg\,min}_{i}J_{ij},i^{-}=\operatorname*{arg\,max}_{i}J_{ij}\}$ (3.12)
and update $\bm{\theta}_{t+1}$ by minimizing the empirical loss of Eq. (3.7) in Line 14. The detailed proof of theorems and lemmas are provided in the Appendix B.
@@ 4.1 Experimental Setup
Data. In our experiments, we use Open Assistant dataset (Köpf et al., 2024) and only reserve about 3.4K human-annotated examples as the seed SFT data $\mathcal{D}_{\text{S}}$. To construct the unlabeled prompt dataset $\mathcal{D}_{\text{U}}$, we mix prompts of $\mathcal{D}_{\text{S}}$ with the train split of each downstream task (only reserve the prompts) including (1) ARC-Easy/Challenge (Clark et al., 2018), (2) OpenBookQA (Mihaylov et al., 2018), (3) SIQA (Sap et al., 2019), and (4) GSM8K (Cobbe et al., 2021). Finally, this process results in a total of 21K prompts in $\mathcal{D}_{\text{U}}$, which we distribute equally across iterative self-rewarding trainings.
@@
Models. Due to limited computational resources, we mainly conduct experiments with two LLMs with about 7B parameters, including Llama-3 (Dubey et al., 2024) and Llama-2 (Touvron et al., 2023), both of which are widely used. Note that the proposed framework is designed for LLMs of any size, while validating our findings on other LLMs will be our future work.
@@
Baseline Methods. We mainly compare our method with SRLM (Yuan et al., 2024) which uses the same LLM to serve as both the policy and reward model to generate preference data for iterative training. Additionally, we introduce a variant of RL with AI feedback (Guo et al., 2024a), referred to as “Oracle”. In this variant, the reward model in SRLM is replaced with an external reward model to demonstrate the upper bound performance of SRLM. Specifically, we use InternLM2 (Cai et al., 2024), a specialized trained reward model, to provide the reward scores for the generated responses. We further enhance Oracle’s rewarding by leveraging the labels from downstream tasks to improve the rewarding accuracy. To compare the regularization method, we introduce “SRLM + KL” variant which adds a simple regularization term $\lambda[\frac{\pi_{\theta}(y|x)}{\pi_{\text{ref}}(y|x)}-\frac{\pi_{\theta}(y^{\prime}|x)}{\pi_{\text{ref}}(y^{\prime}|x)}]^{2}$ to the original DPO loss, where $\lambda\in[0.1,0.2,\cdots,1.0]$ is a hyperparameter. In order to validate the contribution of our automatically determined consistency rate via ranking correlation, we also introduce a variant “CREAM w/o RC” which replaces the ranking correlation (RC) with a manually set value as the consistency rate, where this value is searched in the range of $[0.1,0.3,\cdots,0.9]$.
@@
Implementation Details. In our experiments, we fine-tune the initial model (M0) on the seed SFT data for $3$ epochs with a learning rate of $61e-$ 6, resulting in model M1. Following SRLM approach, we then iteratively fine-tune the model using the preference learning objective two additional iterations, producing models M2 and M3. In the preference training of each iteration, we set $\beta=0.1$ of DPO, and fine-tune the model for $1$ epoch with a learning rate of $61e-$ 6. All training processes use AdamW optimizer (Loshchilov & Hutter, 2019) with a warmup ratio of $0.1$. For the response sampling stage of all SRLM methods, we use a decoding temperature of $0.8$ and generate $N=5$ responses per prompt. For evaluating downstream tasks, we use greedy decoding to generate answers.
@@ 4.2 Main Results
[…]
Model Method Reward Iteration Arc-Easy Arc-Challenge OpenBookQA SIQA GSM8K Average
Llama-3 Initial - M0 86.29 80.37 86.00 68.58 78.01 79.85
SFT - M1 86.78 80.14 86.40 69.50 78.39 80.24
[…]
Llama-2 Initial - M0 61.07 48.98 62.20 50.36 23.65 49.25
SFT - M1 60.44 48.46 63.20 50.77 23.88 49.35
[…]
The main results are shown in Table 1. From these results, we have the following findings:
Figure 3: Arena of CREAM v.s. SRLM and Oracle, judged by GPT-4o.
(1) The Standard SRLM fails to achieve satisfactory performance, particularly with Llama-2 which has weaker foundation performance, which indicates its limitations for 7B-level LLMs. (2) Compared to SRLM, CREAM achieves a significant improvement, showing the advantage of introducing the proposed regularization. (3) SRLM equipped with an oracle reward model (Oracle) achieves the best performance overall. This highlights a critical challenge for all methods with unreliable rewarding. Notably, for Llama-3, CREAM even outperforms Oracle except on SIQA dataset. This superiority underlines the success of the proposed method in mitigating the rewarding bias issue. (4) The consistent improvements of CREAM across iterations show the effectiveness of the proposed regularization method in mitigating the rewarding bias issue. Results of later iterations (up to M6) in Appendix C.1 also demonstrates that adding such regularization can help prevent the model from degeneration in the long term. (5) Though SRLM + KL method improves performance of SRLM via regularization, its overall performance across iterations is still inferior to CREAM, indicating CREAM performs better than directly restricting the KL divergence. (6) CREAM consistently outperforms “CREAM w/o RC” which manually set the consistency rate. And using the dynamically calculated consistency rate can eliminate the costs of hyperparameter search.
Alignment Arena. To more directly compare the alignment performance of CREAM, we further show the Arena win-rate of our method in Figure 3. We can find that CREAM can beat baseline methods with the same iteration, which confirms the superiority of the proposed regularized self-rewarding. Also, the win and tie rates of CREAM increase with the iteration against stronger models such as Oracle and GPT-4o, indicating the consistent performance improvements across iterations.
@@ 4.3.1 Analysis of Rewarding
Rewarding Consistency. We first examine the consistency of rewarding of different methods using their corresponding models from the last iteration in Table 2. Here, we use the proposed Consistency Rate $\mathcal{C}$, Kendall correlation coefficient $\tau$, Spearman correlation coefficient, and TopOrder metrics to measure the consistency, where this metric evaluates whether the final paired preference data remains the same, calculated by: $\text{TopOrder}_{j}=\operatorname{\mathds{1}}\left[\operatorname*{arg\,min}J_{j}=\operatorname*{arg\,min}K_{j}\right]\cdot\operatorname{\mathds{1}}\left[\operatorname*{arg\,max}J_{j}=\operatorname*{arg\,max}K_{j}\right],$ where $J_{j}$ and $K_{j}$ are the rankings provided by current model and the last iteration’s model, respectively. This metric assesses whether both the least preferred and most preferred responses are consistently ranked across iterations. The results confirm that SRLMs exhibit a rewarding consistency issue, indicating that the generated preference data contains noise. In contrast, CREAM keeps the ranking consistency across iterations thanks to the regularized training.
Table 2: Ranking consistency of CREAM and SRLM across iterations using Llama-3.
@@
Prompt Rewarding v.s. DPO Rewarding. As aforementioned, 7B level LLMs struggle with generating accurate rewards when using LLM-as-a-Judge prompting due to their limited capability. Both Figure 5 and Figure 4 clearly show that prompt rewarding is not effective for SRLM with small LLMs, as the performance starts to decrease at the first iteration (M1 $\rightarrow$ M2) when trained on the self-rewarded preference data. In contrast, the adopted DPO rewarding method can be more suitable, since it is intrinsically aligned with the model’s learning objective.
Ranking Accuracy. The ranking accuracy is crucial for self-rewarding methods, as it directly affects the quality of the self-generated preference data for training. Figure 4 provides an intuitive comparison of the rewarding performance. The results include the ranking accuracy on self-generated preference data and the RewardBench (Lambert et al., 2024), both of which is formulated to predict the preferred one between two responses. We use the preference data obtained by self-rewarding with ground truth ranking labels, for testing the model’s in-domain ranking performance. RewardBench is used to assess models’ generalization beyond the training domain. CREAM consistently achieves higher ranking accuracy than baseline methods, which promises more reliable data for training. Even though the overall ranking accuracy ($\leq 70\%$) is not satisfying, the introduced consistency regularization can help mitigate the negative impacts of potentially noisy ranking annotations.